2-3 points Rolff M8 8.7.089 iment with n-60 and ρ-06, use the normal distribution to estimate the following. (See Example 18. Round your answers to four decimal pl (a) = 40) (b) prx-28) (c) PLX = 32) Need Help? Teh to 01-3 points RolFM8 8.7.091. Given the Bernoulli experiment with n-12 and ρ-0.5. use the normal distribution to estimate the following. (See Example 20, Round your answers to four decimal pl (a) P(4 <X<8) Need Help? İİİEl Lemoneel roj 22 points Prevous Answers RorFMB 8.7.097
Answer #1

Q 1)  

The conditions for Normal approximation to binomial distribution is

np > 5 = 60*0.6 = 36 > 5

nq > 5 = 60*0.4 = 24 > 5

Hence we use normal approximation

Mean = n p = 60*0.6 = 36

Standard deviation

σーVnpq = V60 * 0.6 * 0.4 = 3.7947

a)

P(X= 40) =P(39.5<X< 40.5) =P(\frac{39.5 -36}{3.79}<Z<\frac{40.5-36}{3.79})

P(0.92<Z<1.19) = 0.3830 -0.3212 = 0.0618

b)

P(X= 28) =P(27.5<X< 28.5) =P(\frac{27.5 -36}{3.79}<Z<\frac{28.5-36}{3.79})

P(-2.24<Z< -1.98) = 0.4875-0.4761 =0.0114

c)

P(X= 32) =P(31.5<X< 32.5) =P(\frac{31.5 -36}{3.79}<Z<\frac{32.5-36}{3.79})

P(-1.19<Z<-0.92)=0.3830-0.3212 =0.0618